OpenStudy (anonymous):

4. [9.01] Determine whether the graph of y = −5x2 + 3x + 8 opens up or down and whether it has a maximum or minimum point. (5 points) Down; Minimum Down; Maximum Up; Minimum Up; Maximum

OpenStudy (anonymous):

6. [9.02] What are the solutions of 6x2 = 12x? (5 points) x = 0, x = −2 x = 0, x = 2 x = 6, x = 2 x = 6, x = −2 8. [9.02] Solve x2 + 4x − 12 = 0. (5 points) x = 2, x = −6 x = −2, x = −6 x = 3, x = −4 x = −3, x = 4 9. [9.02] What are the x-intercept(s) of the graph of y + 20 = x2 − x? (5 points) (−4, 0) and (−5, 0) (-4, 0) and (5, 0) (4, 0) and (−5, 0) (4, 0) and (5, 0)